A firm with a 14% WACC is evaluating two projects for this year's capital budget. After-tax cash flows, including depreciation, are as follows:

0 1 2 3 4 5
Project M -$21,000 $7,000 $7,000 $7,000 $7,000 $7,000
Project N -$63,000 $19,600 $19,600 $19,600 $19,600 $19,600
  1. Calculate NPV for each project. Do not round intermediate calculations. Round your answers to the nearest cent.

    Project M:    $  

    Project N:    $  

    Calculate IRR for each project. Do not round intermediate calculations. Round your answers to two decimal places.

    Project M:       %

    Project N:       %

    Calculate MIRR for each project. Do not round intermediate calculations. Round your answers to two decimal places.

    Project M:       %

    Project N:       %

    Calculate payback for each project. Do not round intermediate calculations. Round your answers to two decimal places.

    Project M:      years

    Project N:      years

    Calculate discounted payback for each project. Do not round intermediate calculations. Round your answers to two decimal places.

    Project M: years

    Project N:      years
Answer #1

Solution:-

To Calculate NPV of the Project-

To Calculate IRR of the Project-

To Calculate MIRR of project-

MIRR = \frac{Fv of Cash Flows}{Present Value of Cash Outflow}^{\frac{1}{n}} - 1

Project M-

Future Value
Year Deposit Amount Compounding Factor @14% Future Value
1 7000 1.689 11822.72
2 7000 1.482 10370.81
3 7000 1.300 9097.20
4 7000 1.140 7980.00
5 7000 1.000 7000.00
Future Value 46270.73

MIRR = \frac{46,270.73}{21,000}^{\frac{1}{5}} - 1

MIRR = 17.12%

Project N-

Future Value
Year Deposit Amount Compounding Factor @14% Future Value
1 19600 1.689 33103.62
2 19600 1.482 29038.26
3 19600 1.300 25472.16
4 19600 1.140 22344.00
5 19600 1.000 19600.00
Future Value 129558.04

MIRR = \frac{129,558.04}{63,000}^{\frac{1}{5}} - 1

MIRR = 15.51%

To Calculate Payback period of project-

Payback Period = \frac{Initial Investment}{Annual Cash Flows}

Project M-

Payback Period = \frac{21,000}{7,000}

Payback Period = 3 years

Project N-

Payback Period = \frac{63,000}{19,600}

Payback Period = 3.21 years

As per HomeworkLib policy we have to answer only first four parts at once. please ask another part as a separate one.

If you have any query related to question then feel free to ask me in a comment.Thanks. Please rate.